S= (2t 2 + 5t 2 ) i^ + 2t 2 j^ - 5k. what is acceleration at time 2 second

S= (2t+ 5t2) i^ + 2t2 j^ - 5k. what is acceleration at time 2 second 

Grade:12th pass

1 Answers

Arun
4 years ago
Dear Krishna
 
V = dS/dt = 14 t i + 4 t j – 0
hence acceleration a = dV/dt = d2S/dt2 = 14 i + 4j
 
hence after two seconds it will be 14 i + 4j = sqrt (196 + 16) = sqrt (212)